# Stringlet Excitation Model of the Boson Peak

### v2

@@ III Damped dynamics in cooled liquids above the glass transition
which is explicitly demonstrated in the bottom panel of Fig.5. The apparent power law $\zeta$ gradually moves from the Debye value, $\zeta=2$, for small damping towards a linear behavior, $\zeta=1$, for very large damping. As a consequence, for a reasonable value of damping, one expects an apparent power-law in between these two values. A power $\zeta\approx 4/3$ has been observed in superionic UO2 heated crystal Zhang et al. (2019) and a similar power has also been found experimentally in neutron scattering measurements on a Cu46Zr54 metallic glass near room temperature Suck et al. (1980). It is tempting to rationalize such an intermediate power-law frequency scaling of $g(\omega)$ using the argument above and ultimately attribute its origin to the strong stringlet damping emerging at high temperatures. Unfortunately, we are not aware of a systematic simulation study of this effective power-law in a wide range of temperatures, which would be necessary to verify our hypothesis.
@@
$\frac{g_{s}(\omega)}{\omega^{2}}=2\pi\,p_{0}\,v\,\frac{e^{-\frac{\pi\,v}{\lambda\sqrt{\gamma^{2}+\omega^{2}}}}}{\omega\left(\gamma^{2}+\omega^{2}\right)^{3/2}}.$ (21)
@@
This depends explicitly on the value of the stringlet speed and average length. However, we can obtain a universal result by considering a dimensionless stringlet damping $\tilde{\gamma}$ defined as the ratio between the stringlet damping and the BP frequency at $\gamma=0$, or equivalently at zero/low temperature,
@@
$\gamma^{{{*}}}\equiv\left(\frac{\gamma}{\omega^{(0)}_{BP}}\right)^{{{*}}}=0.568\,,$ (26)
@@ IV A simple physical model for a dimensionless stringlet damping
These unstable modes in heated crystalline and glass materials can be calculated from a direct instantaneous normal mode analysis where it is found that these general classes of materials appear rather similar from this perspective Madan and Keyes (1993). Clapa et al. Clapa, Kottos, and Starr (2012) have made the interesting observation that the ratio of the fraction of localized unstable modes to stable localized modes in a model binary Lennard-Jones glass-forming material is rather similar to the ratio of all unstable modes to stable modes estimated from systematic INM analysis of localized and delocalized stable and unstable normal modes over a wide range of temperatures. These findings are relevant to the present analysis since Zhang et al. Zhang et al. (2021a) suggested the identification of the stringlets with localized stable modes in the material and that the the relaxational contribution to fast dynamics arises from the loss of stability of these same modes, where the relaxation process involves irreversible particle exchange motion rather than the reversible periodic motion required of stable modes. String-like collective motion on a ps taking the form of both reversible and reversible motion was directly observed by Zhang et al., supporting this physical interpretation of both the boson peak and the fast relaxation dynamics in terms of stable and unstable modes taking the geometrical expression of string-like displacement motion. It is noted that all modes determined from instantaneous normal mode (INM) analysis are stable at zero temperature and this picture of fast dynamics relaxation simply involves the suggestion that the localized stable modes found at zero temperature bifurcate into stable and unstable localized modes fractions at elevated temperature. This situation is highly plausible because a similar bifurcation of the stable modes in the zero temperature limit into families of stable and unstable modes at finite temperature is a general finding of essentially any INM analysis of condensed materials in their solid and liquid states. Mode damping at elevated temperatures is then a general phenomenon that can be expected to occur for both the localized stringlet modes, which we interpret as the ultimate physical origin of the boson peak, and structural relaxation on a ps timescale (“fast relaxation”), which we correspondingly attribute to unstable striglets.
@@
Supporting this theoretical interpretation of the origin of both the boson peak and fast relaxation in cold condensed materials broadly, it has been repeatedly observed experimentally that a close inter-relationship exists between the “vibrational contribution” of scattering identified with the boson peak and the “relaxational contribution” to scattering intensity in the fast dynamics regime. This has been evidenced by the constancy of ratio of the intensities of depolarized and polarized light scattering, the polarization ratio at fixed temperature, as well as the light to excitation coupling coefficient $C(\omega)$ for both polarized and depolarized light scattering observations exhibiting the same anomalous near linear frequency dependence over a large frequency regime. Notably, this frequency scaling of $C(\omega)$ is consistent with the expectation of 1-dimensional string structures Novikov (1990); Zhang et al. (2021a). As noted previously, the fast dynamics relaxation time obtained from Brillouin scattering measurements has been found to scale inversely with the boson peak frequency Novikov (1997). These and closey related observations regarding the strong interrelationship between the boson peak and relaxational dynamics in the fast dynamics regime are summarized by Ngai et al. in Ref.Ngai, Sokolov, and Steffen (1997). Finally, we point out that only the relative intensities of these elastic and inelastic scattering features, corresponding to the relative contributions of the vibrational and relaxational dynamics of the material, vary with temperature, suggesting a corresponding shift of the fraction of stable to unstable normal modes. Our mathematical model of strong “damping”, introduced below, is entirely consistent with this phenomenology and relates the temperature dependence of the relative fraction of these modes to the temperature dependence of the non-ergodicity parameter, quantifying the height of the intermediate scattering function at the end of the fast beta relaxation process.
@@
This novel model of the fast dynamics of condensed materials relies on two important computational observations. (I) The observed string-like collective motion on a ps timescale completely dominates the magnitude of the mean particle displacement $\langle u^{2}\rangle$ on the fast beta relaxation time because of the relative large displacements of the atoms involved in the stringlet motion Betancourt P., Starr, and Douglas (2018). Notably, $\langle u^{2}\rangle$ does not describe a uniform increase in the amplitude of vibrational motion of the particles in the lattice as you might expect in a harmonic crystal. The anharmonic interparticle interactions of particles at finite temperature make materials rather generally “dynamically heterogeneous” even on a ps timescale, and the string-like particle collective exchange motion is closely tied to this phenomenon. (II) The magnitude of the non-ergodicity parameter of the self-intermediate scattering function on a ps timescale is a uniformly decaying function of $\langle u^{2}\rangle$ in both crystalline and glass-forming materials Pazmiño Betancourt, Douglas, and Starr (2014); Xu, Douglas, and Xu (2023a). This means that the non-ergodicity parameter encodes information about the relative stability of the vibrational modes responsible for the boson peak, and thus the average propensity of these modes to be damped.
@@
It is also helpful for our modeling that recent work has shown that the $T$ dependence of the non-ergodicity parameter can be apparently described by a near-universal function of the mean square particle displacement $\langle u^{2}\rangle$ at the fast beta relaxation time simply because the non-Gaussian parameter is inherently small in the fast dynamics regime of relaxation Betancourt P., Starr, and Douglas (2018); Xu, Douglas, and Xu (2023a). The model just described leads to a precise estimate of $\tilde{\gamma}(T)$ given by,
@@
where $\sigma$ the average interatomic distance estimated from the first peak position of the static structure factor. $T_{o}$ is the temperature at which $\langle u^{2}\rangle$ extrapolates to zero and it indicates the temperature at which the glass-formation process “ends” (see Supplementary Information in Zhang et al. (2021b) for more details). The constant $7.6$ in Eq.(27) derives from the average interparticle distance of the atoms estimated from the peak of the static structure expressed in units of sigma and has been validated using MD simulations in Xu, Douglas, and Xu (2023b). We notice that $\tilde{\gamma}(T)$ ranges between between 00 and $1$, just as the ratio of the fraction of unstable localized and delocalized modes to their stable counterparts. In particular, the fraction of unstable modes is likewise known to increase with temperature as anharmonic interparticle interactions become progressively more prevalent Keyes (1997), and as more normal modes corresponding become progressively unstable, i.e., “damped”. Finally, we note that the expression in Eq.(27) applies quantitatively to polymer materials having a wide range of structures and a wide range of crystalline materials.
@@
Figure 6: Model estimate of the dimensionless stringlet damping $\tilde{\gamma}(T)$ as a function of temperature based on the identification of the intensity in the fast relaxation process as arising the damping of localized stable modes identified as being stringlets in previous simulations Zhang et al. (2021a). In the inset of the bottom panel, the data for the mean squared displacement $\langle u^{2}\rangle$ (in units of $\sigma^{2}$) at the fast beta relaxation time Betancourt P., Starr, and Douglas (2018) in the Al-Sm metallic glass system reproduced from Zhang et al. (2021b). The horizontal dashed line indicates the critical damping value $\tilde{\gamma}^{{{*}}}\approx 0.568$ which corresponds to the characteristic temperature $T^{{{*}}}\approx 848$ K.
@@
Figure 7: The reduced stringlet density of states as a function of the frequency normalized by the zero temperature (or equivalently $\gamma=0$) BP frequency for different temperatures between 00 K to $1350$ K (from blue to red). Solid lines are used to indicate the regimes in which the reduce DOS does not vary much with temperature. The filled black line is the result at zero temperature or $\gamma=0$. The dashed black line is the reduced stringlet DOS at $T^{{{*}}}\approx 848$ K at which the BP simply disappears.
@@
In order to test this model and its predictions we have used the simulation data for the Al-Sm metallic glass system in Zhang et al. (2021a, b) for which the temperature dependence of the mean square displacement is known (see inset in Fig.7). Moreover, we have used the fact that $\sigma\approx 2.8\,$Angstrom at $500$ K for this material and for simplicity we assume $\sigma$ is approximately independent of temperature. We plot our estimate of $\tilde{\gamma}(T)$ in Fig.6. We see that $\tilde{\gamma}(T)$ varies slowly below $\approx 650$ K. It then strongly increases when the temperature becomes larger, between $650$ K and $1200$ K, and then plateaus at very high temperatures. Above $\approx 650$ K, $\langle u^{2}\rangle$ does not vary linearly as for particles subject to harmonic interparticle interaction, so the large and non-linear variation of $\langle u^{2}\rangle$ reflected in the strong increase of $\tilde{\gamma}$ provides clear evidence of large anharmonic interactions in the material where the boson peak is predicted by this model to become significantly damped. In Fig.7, we show the corresponding reduced stringlet density of states as a function of the frequency normalized by the $T=0$, or $\gamma=0$, BP frequency where the dimensionless damping $\tilde{\gamma}$ is modelled based on Eq.(27).
@@
Evidently, the position of the BP approximately does not vary significantly with temperature inside the glass state (in Ref.Zhang et al. (2021a), the glass transition temperature was estimated to be close to $500$ K for the metallic glass under consideration) and it starts to exhibit appreciable softening only above this temperature. As noted above, the BP peak disappears above a characteristic temperature $T^{{{*}}}\approx 848$ K in the AL-Sm glass-forming material, corresponding to the analytically derived critical damping condition discussed in the previous section. Above $T^{{{*}}}$, the stringlet excitations are predicted to be fully overdamped and the DOS then approaches the linear behavior $g_{s}(\omega)\sim\omega$ of a simple fluid. Notice that at elevated temperatures, above $\approx 1200$ K, the VDOS does not vary much as a simple consequence of the dimensionless damping $\tilde{\gamma}$ (Fig.6) approaching a constant value. All these features are in good qualitative agreement with experimental and simulation data. Let us emphasize that the scale at which strong temperature variations appear should relate to the scale at which strong deviations from the scaling $\omega_{BP}\sim G^{1/2}$ are observed, suggesting a possible common origin for these two phenomena.
@@
We conclude with a brief discussion about the critical temperature $T^{{{*}}}\approx 848$. This scale is well above $T_{g}\approx 500$ K, but close to the onset temperature $T_{A}$ defining the onset of Arrhenius dynamics for glassy dynamics which has been independent estimated to be around $900$ K Zhang et al. (2021a). This temperature exactly corresponds to where the intermediate scattering function first develops a discernible multi-step decay, and at which the non-ergodicity parameter becomes measurable. The structural relaxation time typically becomes on the order of a ps at this characteristic temperature and non-Arrhenius relaxation, stretched exponential decay, etc., also first arise below this temperature. The possibility that all these phenomena are just downstream consequences of the stringlets acting on ps timescale is an interesting possibility to consider in the future.
@@ V Outlook
In this work, we showed that a simple theoretical model based on a distribution of 1D vibrating strings Lund (2015) – the stringlet theory of the boson peak – provides an excellent description of the BP frequency $\omega_{BP}$ for 2D and 3D amorphous systems in the glass state, where the damping of these excitations is reasonably be assumed to be negligible. In particular, this theoretical framework gives an analytical, and parameter-free, prediction for $\omega_{BP}$ that is in quantitative agreement with recent simulation data at $T=0$ Hu and Tanaka (2022, 2023). Moreover, the observed scaling of the BP frequency with the shear modulus in this type of zero temperature simulation, $\omega_{BP}\sim G^{1/2}$, follows as an analytic prediction from the stringlet model.
@@
We tentatively explored the stringlet model in the relatively high temperature regime of glass-forming liquids above $T_{g}$ where some damping of the stringlet modes can naturally be expected. In this temperature regime, we addressed the qualitative temperature dependence of the boson peak frequency and its disappearance at high temperatures, in connection with the dynamic correlation length introduced phenomenologically in previous work investigating boson peak measurements Hong, Novikov, and Sokolov (2011a, b). The softening of $\omega_{BP}$ upon heating is reasonably reproduced by the stringlet model and attributed mainly to the change of the stringlet length and transverse sound velocity with temperature. Our qualitative analysis suggests that the stringlet modes become completely overdamped in the liquid regime at sufficiently high temperatures. In association with this phenomenon, we observe the emergence of a density of states that changes to a linear variation with frequency rather than the quadratic law expected for the Debye theory of solids, a transition noted in recent studies attempting to define a density of states appropriate to materials in their liquid state Jin et al. (2023); Keyes (1997); Zaccone and Baggioli (2021); Stamper et al. (2022). We finally mention that recent studies Zhang et al. (2021a); Zhang and Douglas (2013b); Betancourt P., Starr, and Douglas (2018); Hu and Tanaka (2022) of the boson peak frequency as a function of temperature show a nearly linear behavior of $\omega_{BP}$ with $G$, rather than $\propto G^{1/2}$, an observation that offers some clues about the variation of the boson peak in the temperature regime where the average stringlet length, the sound velocity and $\gamma$ are all varying in their own individual ways with temperature. In particular, Zhang et al. Zhang et al. (2021a) indicated that since the increase of $\langle u^{2}\rangle$ and the occurrence of fast beta relaxation first becomes appreciable near the Vogel-Fulcher-Tammann temperature at which the $\alpha$-relaxation time extrapolates to infinity. Stringlet damping should also begin to arise near this same temperature. Rossi et al. Caponi et al. (2011) have observed a deviation of the boson peak frequency from the Debye frequency at temperatures of about $200$ K below $T_{g}$, roughly consistent with the arguments of Zhang et al. that softening of the boson peak occur around $T_{o}$. These and other recent measurements of the pressure dependence of the boson peak Niss et al. (2007); Ahart et al. (2017), indicating a deviation from the simple scaling of the boson peak with the transverse sound velocity, also offer some clues about this more complicated regime for temperatures somewhat below or above $T_{g}$. Further measurements along this line should be helpful in understanding the effect of excitation damping on the boson peak.
@@
Our analysis supports the idea that the stringlets correspond to localized modes, which have long been suggested Zhang et al. (2021a); Hu and Tanaka (2022, 2023) to be responsible for the BP anomaly in amorphous systems,. However, the precise nature and origin of these modes remains an object of continued investigation. This identification raises questions about how these modes might be related to the multipolar four-leaf modes Lerner, Düring, and Bouchbinder (2016); Mizuno, Shiba, and Ikeda (2017) that have recently been discussed as being related to the the emergence of the boson peak. These structures would appear to be much larger than the stringlets suggesting to us the possibility that that stringlets may account for their substructure. This possibility also deserves further investigation.
@@
In order to augment the predictive nature of the stringlet theory, and to develop it into a full-fledged theory, clearly more work needs to be done. First, a theoretical model is required for the first-principle prediction of the stringlet length distribution $p(l)$, without resorting to simulation data. A promising avenue would be to combine Lund model Lund (2015) with the the well-developed thermodynamic theory of equilibrium polymers discussed in Refs. Pazmiño Betancourt, Douglas, and Starr (2014); Douglas, Dudowicz, and Freed (2008, 2006). Second, it would be interesting to directly estimate of the average string length $\lambda$ as a function of temperature, and compare it with independent estimates of the dynamic correlation length based on Eqs.(9)-(10), to verify that that this length can be interrelated as cooperativity length, as proposed heuristically in some experimental studies Hong, Novikov, and Sokolov (2011a, b). Specifically, Eqs.(9)-(10) are equivalent to the “standard” relation taken to define the dynamic correlation length by Hong et al. Hong, Novikov, and Sokolov (2011a, b) and others Kalampounias, Yannopoulos, and Papatheodorou (2006b), when the parameter $S$ of these experimental studies is taken to have the specific predicted value, $S=\pi/(d+1)$.

### v3

@@ III Damped dynamics in cooled liquids above the glass transition
which is explicitly demonstrated in the bottom panel of Fig.5. The apparent power law $\zeta$ gradually moves from the Debye value, $\zeta=2$, for small damping towards a linear behavior, $\zeta=1$, for very large damping. As a consequence, for a reasonable value of damping, one expects an apparent power-law in between these two values. A power $\zeta\approx 4/3$ has been observed in superionic UO2 heated crystal [96] and a similar power has also been found experimentally in neutron scattering measurements on a Cu46Zr54 metallic glass near room temperature [97]. It is tempting to rationalize such an intermediate power-law frequency scaling of $g(\omega)$ using the argument above and ultimately attribute its origin to the strong stringlet damping emerging at high temperatures. Unfortunately, we are not aware of a systematic simulation study of this effective power-law in a wide range of temperatures, which would be necessary to verify our hypothesis.
@@
$\frac{g_{s}(\omega)}{\omega^{2}}=\mathcal{N}_{s}\,\frac{e^{-\frac{\pi\,v}{\lambda\sqrt{\gamma^{2}+\omega^{2}}}}}{\omega\left(\gamma^{2}+\omega^{2}\right)^{3/2}}.$ (23)
@@
This depends explicitly on the value of the stringlet speed and average length $\langle l\rangle=\lambda+1$. However, we can obtain a universal result by considering a dimensionless stringlet damping $\tilde{\gamma}$ defined as the ratio between the stringlet damping and the BP frequency at $\gamma=0$, or equivalently at zero/low temperature,
@@
$\tilde{\gamma}^{{{*}}}\equiv\left(\frac{\gamma}{\omega^{(0)}_{BP}}\right)^{{{*}}}=0.568\,,$ (28)
@@ IV A simple physical model for a dimensionless stringlet damping
These unstable modes in heated crystalline and glass materials can be calculated from a direct instantaneous normal mode analysis where it is found that these general classes of materials appear rather similar from this perspective [98]. Clapa et al. [99] have made the interesting observation that the ratio of the fraction of localized unstable modes to stable localized modes in a model binary Lennard-Jones glass-forming material is rather similar to the ratio of all unstable modes to stable modes estimated from systematic INM analysis of localized and delocalized stable and unstable normal modes over a wide range of temperatures. These findings are relevant to the present analysis since Zhang et al. [32] suggested the identification of the stringlets with localized stable modes in the material and that the the relaxational contribution to fast dynamics arises from the loss of stability of these same modes, where the relaxation process involves irreversible particle exchange motion rather than the reversible periodic motion required of stable modes. String-like collective motion on a ps taking the form of both reversible and reversible motion was directly observed by Zhang et al., supporting this physical interpretation of both the boson peak and the fast relaxation dynamics in terms of stable and unstable modes taking the geometrical expression of string-like displacement motion. It is noted that all modes determined from an instantaneous normal mode (INM) analysis are stable at zero temperature and this picture of fast dynamics relaxation simply involves the suggestion that the localized stable modes found at zero temperature bifurcate into stable and unstable localized modes fractions at elevated temperature. This situation is highly plausible because a similar bifurcation of the stable modes in the zero temperature limit into families of stable and unstable modes at finite temperature is a general finding of essentially any INM analysis of condensed materials in their solid and liquid states. Mode damping at elevated temperatures is then a general phenomenon that can be expected to occur for both the localized stringlet modes, which we interpret as the ultimate physical origin of the boson peak, and structural relaxation on a ps timescale (“fast relaxation”), which we correspondingly attribute to unstable striglets.
@@
Supporting this theoretical interpretation of the origin of both the boson peak and fast relaxation in cold condensed materials broadly, it has been repeatedly observed experimentally that a close inter-relationship exists between the “vibrational contribution” of scattering identified with the boson peak and the “relaxational contribution” to scattering intensity in the fast dynamics regime. This has been evidenced by the constancy of ratio of the intensities of depolarized and polarized light scattering, the polarization ratio at fixed temperature, as well as the light to excitation coupling coefficient $C(\omega)$ for both polarized and depolarized light scattering observations exhibiting the same anomalous near linear frequency dependence over a large frequency regime. Notably, this frequency scaling of $C(\omega)$ is consistent with the expectation of 1-dimensional string structures [44; 32]. As noted previously, the fast dynamics relaxation time obtained from Brillouin scattering measurements has been found to scale inversely with the boson peak frequency [100]. These and closey related observations regarding the strong interrelationship between the boson peak and relaxational dynamics in the fast dynamics regime are summarized by Ngai et al. in Ref.[101]. Finally, we point out that only the relative intensities of these elastic and inelastic scattering features, corresponding to the relative contributions of the vibrational and relaxational dynamics of the material, vary with temperature, suggesting a corresponding shift in the relative fraction of stable to unstable normal modes. Our mathematical model of strong “damping”, introduced below, is entirely consistent with this phenomenology and relates the temperature dependence of the relative fraction of these modes to the temperature dependence of the non-ergodicity parameter, quantifying the height of the intermediate scattering function at the end of the fast beta relaxation process.
@@
This novel model of the fast dynamics of condensed materials relies on two important computational observations. (I) The observed string-like collective motion on a ps timescale completely dominates the magnitude of the mean particle displacement $\langle u^{2}\rangle$ on the fast beta relaxation time because of the relative large displacements of the atoms involved in the stringlet motion [37]. Notably, $\langle u^{2}\rangle$ does not describe a uniform increase in the amplitude of vibrational motion of the particles in the lattice as you might expect in a harmonic crystal. The anharmonic interparticle interactions of particles at finite temperature make materials rather generally “dynamically heterogeneous” even on a ps timescale, and the string-like particle collective exchange motion is closely tied to this phenomenon. (II) The magnitude of the non-ergodicity parameter of the self-intermediate scattering function on a ps timescale is a uniformly decaying function of $\langle u^{2}\rangle$ in both crystalline and glass-forming materials [59; 102]. This means that the non-ergodicity parameter encodes information about the relative stability of the vibrational modes responsible for the boson peak, and thus the average propensity of these modes to be damped.
@@
It is also helpful for our modeling that recent work has shown that the $T$ dependence of the non-ergodicity parameter can be apparently described by a near-universal function of the mean square particle displacement $\langle u^{2}\rangle$ at the fast beta relaxation time simply because the non-Gaussian parameter is inherently small in the fast dynamics regime of relaxation [37; 102]. The model just described leads to a precise estimate of $\tilde{\gamma}(T)$ given by,
@@
where $\sigma$ the average interatomic distance estimated from the first peak position of the static structure factor. $T_{o}$ is the temperature at which $\langle u^{2}\rangle$ extrapolates to zero, indicating the temperature at which the glass-formation process “ends” (see Supplementary Information in [81] for more details). The constant $7.6$ in Eq.(29) derives from the average interparticle distance of the atoms estimated from the peak of the static structure expressed in units of sigma and has been validated using MD simulations in [103]. We notice that $\tilde{\gamma}(T)$ ranges between between 00 and $1$, just as the ratio of the fraction of unstable localized and delocalized modes to their stable counterparts. In particular, the fraction of unstable modes is likewise known to increase with temperature as anharmonic interparticle interactions become progressively more prevalent [86], and as more normal modes corresponding become progressively unstable, i.e., “damped”. Finally, we note that the expression in Eq.(29) applies quantitatively to polymer materials having a wide range of structures and a wide range of crystalline materials.
@@
Figure 6: Model estimate of the dimensionless stringlet damping $\tilde{\gamma}(T)$ as a function of temperature based on the identification of the intensity in the fast relaxation process as arising the damping of localized stable modes identified as being stringlets in previous simulations [32]. The horizontal dashed line indicates the critical damping value $\tilde{\gamma}^{{{*}}}\approx 0.568$ which corresponds to the characteristic temperature $T^{{{*}}}\approx 848$ K.
@@
Figure 7: The reduced stringlet density of states as a function of the frequency normalized by the zero temperature (or equivalently $\gamma=0$) BP frequency for different temperatures between 00 K to $1350$ K (from blue to red). Solid lines are used to indicate the regimes in which the reduce DOS does not vary much with temperature. The filled black line is the result at zero temperature or $\gamma=0$. The dashed black line is the reduced stringlet DOS at $T^{{{*}}}\approx 848$ K at which the BP simply disappears. In the inset, the data for the mean squared displacement $\langle u^{2}\rangle$ (in units of $\sigma^{2}$) at the fast beta relaxation time [37] data for the Al-Sm metallic glass system are taken from [81].
@@
In order to test this model and its predictions we have used the simulation data for the Al-Sm metallic glass system in [32; 81] for which the temperature dependence of the mean square displacement is known (see inset in Fig.7). Moreover, we have used the fact that $\sigma\approx 2.8\,$Angstrom at $500$ K for this material and for simplicity we assume $\sigma$ is approximately independent of temperature. We plot our estimate of $\tilde{\gamma}(T)$ in Fig.6. We see that $\tilde{\gamma}(T)$ varies slowly below $\approx 650$ K. It then strongly increases when the temperature becomes larger, between $650$ K and $1200$ K, and then plateaus at very high temperatures. Above $\approx 650$ K, $\langle u^{2}\rangle$ does not vary linearly as for particles subject to harmonic interparticle interaction, so the large and non-linear variation of $\langle u^{2}\rangle$ reflected in the strong increase of $\tilde{\gamma}$ provides clear evidence of large anharmonic interactions in the material where the boson peak is predicted by this model to become significantly damped. In Fig.7, we show the corresponding reduced stringlet density of states as a function of the frequency normalized by the $T=0$, or $\gamma=0$, BP frequency where the dimensionless damping $\tilde{\gamma}$ is modelled based on Eq.(29).
@@
Evidently, the position of the BP approximately does not vary significantly with temperature inside the glass state (in Ref.[32], the glass transition temperature was estimated to be close to $500$ K for the metallic glass under consideration) and it starts to exhibit appreciable softening only above this low temperature onset. As noted above, the BP peak disappears above a characteristic temperature $T^{{{*}}}\approx 848$ K in the AL-Sm glass-forming material, corresponding to the analytically derived critical damping condition discussed in the previous section. Above $T^{{{*}}}$, the stringlet excitations are predicted to be fully overdamped and the DOS then approaches the linear behavior $g_{s}(\omega)\sim\omega$ of a simple fluid. Notice that at elevated temperatures, above $\approx 1200$ K, the VDOS does not vary much as a simple consequence of the dimensionless damping $\tilde{\gamma}$ (Fig.6) approaching a constant value. All these features are in good qualitative agreement with experimental and simulation data. Let us emphasize that the scale at which strong temperature variations appear should relate to the scale at which strong deviations from the scaling $\omega_{BP}\sim G^{1/2}$ are observed, suggesting a possible common origin for these two phenomena.
@@
We conclude with a brief discussion about the critical temperature $T^{{{*}}}\approx 848$. This scale is well above $T_{g}\approx 500$ K, but close to the onset temperature $T_{A}$ defining the onset of Arrhenius dynamics for glassy dynamics which has been independent estimated to be around $900$ K [32]. This temperature exactly corresponds to where the intermediate scattering function first develops a discernible multi-step decay, and at which the non-ergodicity parameter becomes measurable. The structural relaxation time typically becomes on the order of a ps at this characteristic temperature and non-Arrhenius relaxation, stretched exponential decay, etc., also first arise below this temperature. The possibility that all these phenomena are just downstream consequences of the stringlets acting on ps timescale is an interesting possibility to consider in the future.
@@ V Outlook
In this work, we showed that a simple theoretical model based on a distribution of 1D vibrating strings [55] – the stringlet model of the boson peak – provides a good description of the BP frequency $\omega_{BP}$ for 2D and 3D amorphous systems in the glass state, where the damping of these excitations is reasonably be assumed to be negligible. In particular, this theoretical framework gives an analytical, and parameter-free, prediction for $\omega_{BP}$ that is in quantitative agreement with recent simulation data at $T=0$ [35; 36]. Moreover, the observed scaling of the BP frequency with the shear modulus in this type of zero temperature simulation, $\omega_{BP}\sim G^{1/2}$, follows as an analytic prediction from the stringlet model.
@@
We tentatively explored the stringlet model in the relatively high temperature regime of glass-forming liquids above $T_{g}$ where some damping of the stringlet modes can naturally be expected. In this temperature regime, we addressed the qualitative temperature dependence of the boson peak frequency and its disappearance at high temperatures, in connection with the dynamic correlation length introduced phenomenologically in previous work investigating boson peak measurements [60; 61]. The softening of $\omega_{BP}$ upon heating is reasonably reproduced by the stringlet model and attributed mainly to the change of the stringlet length and transverse sound velocity with temperature. Our qualitative analysis suggests that the stringlet modes become completely overdamped in the liquid regime at sufficiently high temperatures. In association with this phenomenon, we observe the emergence of a density of states that changes to a linear variation with frequency rather than the quadratic law expected for the Debye theory of solids, a transition noted in recent studies attempting to define a density of states appropriate to materials in their liquid state [85; 86; 87; 84]. We finally mention that recent studies [32; 34; 37; 35] of the boson peak frequency as a function of temperature show a nearly linear behavior of $\omega_{BP}$ with $G$, rather than $\propto G^{1/2}$, an observation that offers some clues about the variation of the boson peak in the temperature regime where the average stringlet length, the sound velocity and $\gamma$ are all varying in their own individual ways with temperature. In particular, Zhang et al. [32] indicated that since the increase of $\langle u^{2}\rangle$ and the occurrence of fast beta relaxation first becomes appreciable near the Vogel-Fulcher-Tammann temperature at which the $\alpha$-relaxation time extrapolates to infinity. Stringlet damping should also begin to arise near this same temperature. Rossi et al. [104] have observed a deviation of the boson peak frequency from the Debye frequency at temperatures of about $200$ K below $T_{g}$, roughly consistent with the arguments of Zhang et al. that softening of the boson peak occur around $T_{o}$. These and other recent measurements of the pressure dependence of the boson peak [105; 106], indicating a deviation from the simple scaling of the boson peak with the transverse sound velocity, offer some clues about this more complicated regime for temperatures somewhat below or above $T_{g}$. Further measurements along this line should be helpful in understanding the effect of excitation damping on the boson peak.
@@
Our analysis supports the idea that the stringlets correspond to localized modes, which have long been suggested [32; 35; 36] to be responsible for the BP anomaly in amorphous systems,. However, the precise nature and origin of these modes remains an object of continued investigation. This identification raises questions about how these modes might be related to the multipolar four-leaf modes [107; 108] that have recently been discussed as being related to the the emergence of the boson peak. These structures would appear to be much larger than the stringlets suggesting to us the possibility that that stringlets may account for their substructure. This possibility also deserves further investigation.
@@
In order to augment the predictive nature of the stringlet model, and to develop it into a full-fledged theory, clearly more work needs to be done. First, a theoretical model is required for the first-principle prediction of the stringlet length distribution $p(l)$, and in particular the total stringlet number proportional to $\mathcal{N}_{s}$, without resorting to simulation data. A promising avenue would be to combine Lund model [55] with the the well-developed thermodynamic theory of equilibrium polymers discussed in Refs. [59; 109; 110]. Second, it would be interesting to directly estimate of the average string length $\langle l\rangle$ as a function of temperature, and compare it with independent estimates of the dynamic correlation length based on Eqs.(11)-(12), to verify that that this length can be interrelated as cooperativity length, as proposed heuristically in some experimental studies [60; 61]. Specifically, Eqs.(11)-(12) are equivalent to the “standard” relation taken to define the dynamic correlation length by Hong et al. [60; 61] and others [111], when the parameter $S$ of these experimental studies is taken to have the specific predicted value, $S=\pi/(d+1)$, with $d$ the number of spatial dimensions.